### v1

# Lumos-1: On Autoregressive Video Generation from a Unified Model Perspective

Abstract
Autoregressive large language models (LLMs) have unified a vast range of language tasks, inspiring preliminary efforts in autoregressive video generation. Existing autoregressive video generators either diverge from standard LLM architectures, depend on bulky external text encoders, or incur prohibitive latency due to next-token decoding. In this paper, we introduce Lumos-1, an autoregressive video generator that retains the LLM architecture with minimal architectural modifications. To inject spatiotemporal correlations in LLMs, we identify the efficacy of incorporating 3D RoPE and diagnose its imbalanced frequency spectrum ranges. Therefore, we propose MM-RoPE, a RoPE scheme that preserves the original textual RoPE while providing comprehensive frequency spectra and scaled 3D positions for modeling multimodal spatiotemporal data. Moreover, Lumos-1 resorts to a token dependency strategy that obeys intra-frame bidirectionality and inter-frame temporal causality. Based on this dependency strategy, we identify the issue of frame-wise loss imbalance caused by spatial information redundancy and solve it by proposing Autoregressive Discrete Diffusion Forcing (AR-DF). AR-DF introduces temporal tube masking during training with a compatible inference-time masking policy to avoid quality degradation. By using memory-efficient training techniques, we pre-train Lumos-1 on only 48 GPUs, achieving performance comparable to EMU3 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V. Code and models are available at https://github.com/alibaba-damo-academy/Lumos.
1 Introduction
Figure 1: Visualization of examples generated by Lumos-1. Lumos-1 supports text-to-image, image-to-video and text-to-video tasks.
Autoregressive models have demonstrated significant advancement in the field of language processing [6, 44, 70, 45, 3, 37] by unifying diverse language tasks into one single framework and expanding the scale of large language models to unprecedented sizes. Following this trail of research, significant efforts have emerged in autoregressive visual generation [33, 7, 8, 42]. By utilizing similar architectural and generation designs as those used in LLMs, there is substantial potential for advancing LLMs towards a unified model capable of both visual generation and understanding [77, 81, 95, 68].
Pioneering research efforts have been dedicated to instantiating autoregressive visual generation, with a focus on key aspects, spanning improving autoregressive paradigm [46, 67, 8, 42], improving tokenizer format [33] (e.g., from discrete to continuous), and enhancing tokenizer capabilities [64, 1, 74, 58]. However, an autoregressive video generation paradigm that is fully compatible with unified models remains underexplored. Preliminary attempts either exhibit architectural distinctions with LLM architectures [14, 72], rely on external text encoders [14], or trail in generation efficiency [79], leaving much room for advancement.
In this work, we propose Lumos-1, a model that leverages the exact architecture of LLM with minimal architectural modifications to achieve autoregressive video generation, in such a way that it enables video generation without a pre-trained external text encoder (figure 1 visualizes examples generated by Lumos-1). To ensure inference efficiency, we build Lumos-1 upon the discrete diffusion paradigm [81, 63], whose lower bound is the cross-entropy loss employed in MaskGIT [8].
Distinct from texts, videos inherently exhibit strong spatiotemporal correlation. Although previous works have incorporated this prior into generative models by overturning the next-token prediction paradigm [67, 8], we dive into a basic technique that’s widely used in LLMs—Rotary Position Embeddings (RoPE), which is heavily underexplored in autoregressive video generation. Preliminary experiments conducted with Lumos-1 reveal that 1D RoPE is far from optimal, and the incorporation of 3D RoPE facilitates autoregressive generative learning. However, existing 3D RoPE suffers from imbalanced frequency spectrum ranges for temporal and spatial modeling. Building on this insight, we propose MM-RoPE, an improved version of RoPE that better accommodates the structure of video data. This enhancement involves more distributed frequency allocations for more comprehensive context modeling and the appropriate scaling of 3D position information for modality balancing. Though MM-RoPE modifies RoPE for visual tokens, the RoPE for text tokens remains consistent with the one from original LLMs, therefore preserving the language learning capabilities.
Distinct from images, videos inherently exhibit 1) temporal causality between frames and 2) spatial information redundancy between frames. The temporal causality renders the paradigm of autoregressive image generation suboptimal for videos, which relies on global sequence bidirectional dependency [8] or raster-scan order dependency [62]. In response, Lumos-1 resorts to a token dependency strategy that emphasizes intra-frame bidirectional dependencies and inter-frame temporally causal dependencies, which resonates with the nature of video data. Meanwhile, spatial information redundancy degrades the efficacy of training with random mask prediction, as spatial information leakage allows the model to easily predict the masked token by attending to unmasked ones in previous frames. To alleviate this, Lumos-1 introduces Autoregressive Discrete Diffusion Forcing (AR-DF). The core of AR-DF training involves temporal tube masking [69], which repeats a frame-level random mask pattern across the temporal dimension. To some extent, such a design mitigates the issue of frame-wise imbalanced loss [79], thereby enabling more effective learning. The core of AR-DF inference includes an inference-time masking strategy that aligns with partial observation of history during training, enabling inference without degraded frame quality and motion.
We leverage Llama [71] as our architecture and leverage a unified discrete codebook for language and visual data. Through a stage-wise training with GPU memory-friendly techniques, we pre-tarin Lumos-1 from scratch on 60 million images and 10 million videos using only 48 GPUs, achieving performance comparable to EMU3 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V.
2 Related Work
Autoregressive image generation. Spurred by the progress of LLMs, visual generation research is rapidly shifting from diffusion paradigms [55, 24, 93, 15] to an autoregressive one [62, 67, 86, 27, 52, 96, 90, 23, 43, 75, 53], due to its potential of being integrated into unified models [65, 97] with minimal modification. Preliminary research, such as Parti [87], DALL-E [51], MaskGiT [8] and LLamaGen [62], have demonstrated the efficacy of autoregressive generation using discrete tokenizers. MAR [33] incorporates a diffusion-style objective into autoregressive training, thereby accommodating continuous tokens. VAR [67] and FAR [86] reformulate the autoregressive target as next-scale and next-frequency prediction, respectively, improving generation fidelity. Unlike them, we dive into the design of the RoPE technique in visual generation.
Autoregressive video generation. Autoregressive video synthesis follows a trajectory similar to images, but the substantial computational overhead encouraged researchers to explore different granularities of autoregression. At the macro level, methods such as CausVid [84], Pyramidal Flow [29], and MAGI-1 [56] generate video clips by recursively invoking diffusion models. A second line of work adopts a hybrid AR-diffusion strategy that couples a pre-trained diffusion backbone with an external autoregressive planner, as exemplified by ARLON [34] and Mardini [40]. Our focus, however, is on micro AR approaches [21, 96, 88, 89, 72, 30, 77, 14, 79, 85], which treat the entire spatiotemporal token sequence as a single context and generate videos end-to-end with one autoregressive transformer (e.g., Phenaki [72], VideoPoet [30], Loong [79], EMU3 [77], NOVA [14]). Building on this paradigm, we introduce AR-DF, which is compatible with intra-frame bidirectionality and inter-frame temporal causality, enabling effective training and inference with LLM-style architectures.
RoPE for vision-language data. RoPE was first proven effective for LLMs [61, 70] and later adopted in DiTs [47] to inject spatial priors. However, its potential for autoregressive video generation remains heavily underexplored. Related works, including M-RoPE [76], RoPE-2D [2], TAD-RoPE [17] and VideoRoPE [80], incorporated structure priors (i.e., spatiotemporal information or its subsets) into RoPE for better vision-language understanding. VoPE [10] improves high-resolution image generation by continuous resolution learning. We propose MM-RoPE, a distributed and scaled RoPE technique that improves AR video synthesis while remaining plug-and-play for unified multimodal models.
3 Lumos-1
This section aims to introduce the design philosophy behind Lumos-1. In Sec. 3.1, we introduce MM-RoPE that enables better spatiotemporal awareness of an LLM for modeling visual data. In Sec. 3.2, we introduce AR-DF that enables effective training and inference. In Sec. 3.3, we introduce important techniques for implementing Lumos-1, including architecture, memory-friendly techniques, etc.
3.1 Spatiotemporal Correlation Injection via MM-RoPE
Figure 2: (a) Initial exploration of 3D RoPE in autoregressive video generation. The sequence is comprised of text tokens and visual tokens (default arrangement in this paper), with global indices starting from 0. Two schemes are exemplified using one text token and one visual token. In Scheme 1 and 2, temporal, height and width positions start from 0. (b) Details of MM-RoPE compared to M-RoPE. The figure illustrates the distributed channel allocation of MM-RoPE. Temporal, height and width positions are indexed after text tokens.
Preliminaries of 3D RoPE. One de facto design of contemporary LLM is the RoPE [61] technique, whose overall aim is to encode the absolute position with a rotation matrix while incorporating the explicit relative position dependency in the attention mechanism. This can be formulated as:
$\langle f_{q}(\bm{x}_{m},m),f_{k}(\bm{x}_{n},n)\rangle=g(\bm{x}_{m},\bm{x}_{n},m-n)$ (1)
where $f_{q}(\bm{x}_{m},m)$ encodes the position $m$ for the embedding $\bm{x}_{m}$ to obtain the query feature ($f_{k}(\bm{x}_{n},n)$ is analogously defined); $g(\bm{x}_{m},\bm{x}_{n},m-n)$ is a function that defines the inner product between the query and key vectors that explicitly encodes the relative position $m-n$. The resultant form of function $f_{\{q,k\}}$ can be formulated as:
$f_{\{q,k\}}(\bm{x}_{m},m)=\bm{R}^{d}_{\Theta,m}\bm{W}_{q,k}\bm{x}_{m}$ (2)
where $\bm{W}_{q,k}$ is the projection matrix; $\bm{R}^{d}_{\Theta,m}$ is the rotary matrix with pre-defined parameters $\Theta=\{\theta_{i}=\beta^{-2(i-1)/d},i=[1,2,...,d/2]\}$ with $d$ acting as the dimension of the embeddings and $\beta$ acting as the base frequency. In such a formulation, the attention calculation can be rewritten as:
$\begin{split}f_{q}(\bm{x}_{m},m)^{\rm T}f_{k}(\bm{x}_{n},n)&=\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}\bm{R}^{d}_{\Theta,\tau}\bm{W}_{k}\bm{x}_{n}\\ \tau&=n-m\end{split}$ (3)
where the detailed formulation of $\bm{R}^{d}_{\Theta,\tau}$ can be formulated using a base rotary matrix $R_{\theta,\tau}$, with $\theta$ as the frequency and $\tau$ as the relative position:
$\begin{split}\bm{R}^{d}_{\Theta,\tau}&=\begin{bmatrix}R_{\theta_{1},\tau}&0&\cdots&0\\ 0&R_{\theta_{2},\tau}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\cdots&R_{\theta_{d/2},\tau}\\ \end{bmatrix}\\ R_{\theta,\tau}&=\begin{bmatrix}\cos\tau\theta&-\sin\tau\theta\\ \sin\tau\theta&\cos\tau\theta\end{bmatrix}\end{split}$ (4)
However, the application of the original RoPE to modeling visual data remains suboptimal considering the spatiotemporal correlation of visual tokens. One popular type of generative models, diffusion models [25, 59, 24], have improved upon this technique by proposing 3D RoPE that jointly injects spatiotemporal latent coordinates during attention calculation and demonstrates its effectiveness [26, 31, 66]. If we slightly abuse the annotation by denoting $\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}$ and $\bm{W}_{k}\bm{x}_{n}$ as $\bm{X}_{m}^{\rm T}$ and $\bm{X}_{n}$, we can formulate the attention calculation in equation 1 based on 3D RoPE as:
$\bm{X}_{m,t_{s}:t_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{t_{s}+1},\tau_{t}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{t_{e}},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,t_{s}:t_{e}}+$ (5)
$\bm{X}_{m,h_{s}:h_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{h_{s}+1},\tau_{h}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{h_{e}},\tau_{h}}\\ \end{bmatrix}\bm{X}_{n,h_{s}:h_{e}}+$
$\bm{X}_{m,w_{s}:w_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{w_{s}+1},\tau_{w}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{w_{e}},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,w_{s}:w_{e}}$
where $\{t_{s},t_{e}\}=\{0,\frac{2}{16}d\}$, $\{h_{s},h_{e}\}=\{\frac{2}{16}d,\frac{5}{16}d\}$ and $\{w_{s},w_{e}\}=\{\frac{5}{16}d,\frac{1}{2}d\}$ denote the start and end dimension index for encoding temporal, height and width relative position; $\bm{X}_{m,t_{s}:t_{e}}^{\rm T}$ denotes the submatrix extracted from $\bm{X}_{m}^{\rm T}$ using row indices $[2t_{s},2t_{e})$; other matrices are similarly defined.
Figure 3: (a) Validation loss curve of a toy experiment exploring the necessity of 3D RoPE using the 0.5B model; (b) Frequency allocation in the vanilla 3D RoPE; (c) Comparison of rotary speed of the first dimension for temporal and height dimensions (high and middle frequencies) in the vanilla 3D RoPE.
Initial exploration of incorporating 3D RoPE. Our preliminary exploration starts with the introduction of 3D RoPE to autoregressive video generation. We follow previous works [16, 33] to utilize the validation loss to observe the effectiveness due to its strong correlation with evaluation metrics. By default, we use the cross-entropy loss (C-Loss), following the standard LLM training objective. We compare the vanilla LLM RoPE with three schemes, as shown in figure 2: 1) Scheme 1, which allocates the first $1/2$ channels to encode the global position (i.e., the index in the global sequence) and the second $1/2$ channels to encode the temporal, height and width positions with a ratio of $2:3:3$. For text tokens, we only use the first half channels to encode the global positions to ensure the language modeling capability, while for visual tokens, we only use the second half to encode 3D positions; 2) Scheme 2, which extends upon scheme 1 by leveraging the first half channels of the visual tokens to encode the global positions. 3) M-RoPE [76], which uses all channels of the visual tokens to encode the 3D positions. From figure 3(a), we can observe that: 1) The incorporation of the spatiotemporal correlation in RoPE can significantly improve the performance of the model fitting data by comparing vanilla RoPE and Scheme 1. 2) The injection of the raster-scan order position information to visual tokens (i.e., the global position in Scheme 2 of figure 2(a)) can degrade the performance. 3) A comprehensive channel utilization (M-RoPE) is better than partial channel utilization (Scheme 1). Therefore, it is promising to inject such priors in this generative model.
Peering into 3D RoPE and identifying its limitations. Although 3D RoPE proves effective in practice, its design remains suboptimal. In figure 3(b), we visualize how the frequencies are allocated to model temporal, height, and width dimensions. We observe that the temporal channels dominate large frequency spectrum, whereas the height and width channels are relegated to near-zero frequencies. For the sine function, the relative positions $\tau$ (when $\tau>=0$) should not exceed one period to avoid ambiguity since radians beyond $2\pi$ start repeating patterns in the function. Beyond this range, the model cannot distinguish fine-grained positional differences since rotary embeddings of this channel lose uniqueness. For the low-indexed channels, the embeddings can rotate significantly faster than high-indexed channels (as shown in figure 3(c)), leading to accelerated aliasing and loss of embedding uniqueness. For the high-indexed channels, the embeddings rotate so slowly that they lack sufficient resolution to model subtle local changes. Finally, while height and width are symmetrically important, they occupy disproportionately small and different segments of the overall frequency spectrum, reducing their capacity to capture spatial details effectively.
MM-RoPE: a distributed and scaled 3D RoPE mechanism. To elegantly solve the aforementioned limitation, we propose MM-RoPE, a distributed 3D RoPE mechanism. Compared with M-RoPE [76] that is widely adopted in vision-language models, one core idea of MM-RoPE is to encode relative positions across comprehensive frequency spectra for all 3D information. As illustrated in figure 2(b), the RoPE for text tokens in MM-RoPE follows the standard design of an LLM, whereas the RoPE for visual tokens is comprised of multiple meta MM-RoPE components. Within each meta MM-RoPE, we keep the ratio of 3D information identical to 3D RoPE (i.e., $2:3:3$), while minimizing the overall dimensionality to maintain a more distributed design. Concretely, we first allocate channels for temporal information, then symmetrically interleave height and width channels to model spatial information. We can formulate the attention calculation of the first meta MM-RoPE as:
$\bm{X}_{m,0:2}^{\rm T}\begin{bmatrix}R_{\theta_{1},\tau_{t}}&0\\ 0&R_{\theta_{2},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,0:2}+$ (6)
$\bm{X}_{m,2:4}^{\rm T}\begin{bmatrix}R_{\theta_{3},\tau_{h}}&0\\ 0&R_{\theta_{4},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,2:4}+$
$\bm{X}_{m,4:6}^{\rm T}\begin{bmatrix}R_{\theta_{5},\tau_{h}}&0\\ 0&R_{\theta_{6},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,4:6}+$
$\bm{X}_{m,6:8}^{\rm T}\begin{bmatrix}R_{\theta_{7},\tau_{h}}&0\\ 0&R_{\theta_{8},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,6:8}$
where each meta MM-RoPE component comprises $16$ channels; additional components are defined analogously and collectively form the RoPE strategy for visual tokens.
Algorithm 1 AR-DF Training Procedure
Require: video dataset $\mathcal{D}$, training mask ratio $\rho_{tra}$, generative model $G_{\phi}$, number of latent frames $T$
1:for each video $\bm{X}\in\mathcal{D}$ do
[…]
3:    Sample a mask pattern for all frames:
4:     $\rho\sim\mathrm{Uniform}([\rho_{tra},1])$;
5:     $\bm{M}\sim\mathrm{Bernoulli}(1-\rho)$
6:    for $t=1\to T$ do $\triangleright$ Apply $\bm{M}$ to all frames
7:       $\widetilde{\bm{X}}_{v}^{(t)}=\bm{M}\odot\bm{X}_{v}^{(t)}+(1-\bm{M})\odot[\text{MASK}]$
8:    end for
9:    Form $\widetilde{\bm{X}}=\{\bm{X}_{p},\,\widetilde{\bm{X}}_{v}^{(1)},\dots,\widetilde{\bm{X}}_{v}^{(T)}\}$
10:    Generate attention mask $AttnMask$ for $\widetilde{\bm{X}}$
11:    $\widehat{\bm{X}}\leftarrow G_{\phi}(\widetilde{\bm{X}},AttnMask)$
12:    Compute loss $\mathcal{L}(\widehat{\bm{X}},\bm{X},\bm{M})$
13:    Backprop & update $\phi$
14:end for
Algorithm 2 AR-DF Inference Procedure
Require: text prompt $\bm{X}_{p}$, trained model $G_{\phi}$, inference mask ratio $\rho_{inf}$, number of latent frames $T$, number of generation steps $N_{steps}$, number of tokens in a latent frame $N_{f}$, KV cache $\mathcal{C}=\varnothing$, generated latent list $\mathcal{V}=\varnothing$
1:Initialize text cache:
2: Generate text causal mask $AttnMask^{(p)}$ for $\bm{X}_{p}$
3:$\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$triangleright Store cache for the prompt
4:Sample cache mask:
5:$\bm{M}_{inf}\sim\mathrm{Bernoulli}(1-\rho_{inf})$
6:for $t=1\to T$ do $\triangleright$ Reused for all frames
[…]
8:    Generate temporal causal mask $AttnMask^{(t)}$ for $\{\bm{X}_{p},\bm{X}_{v}^{(t)}\}$
9:    Generate $t$-th frame $\bm{X}_{v}^{(t)}$ and append it to $\mathcal{V}$
10:    Cache partial observation of the generated frame:
11:     $\widetilde{\bm{X}}_{v}^{(t)}=\bm{M}_{inf}\odot\bm{X}_{v}^{(t)}+(1-\bm{M}_{inf})\odot[\text{MASK}]$
12:     $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$triangleright Store cache for the $t$-th frame
13:end for
Figure 4: Left: Training algorithm with AR-DF. For simplicity, this only includes training on videos. Right: Inference algorithm with AR-DF. A more detailed version for inference is included in the Appendix.
Figure 5: (a) Details of scaled 3D positions in MM-RoPE. (b) Temporal causal mask used in Lumos-1.
Moreover, for a model that jointly processes text and visual tokens, the interplay between two modalities is significant to ensure the vision-language alignment. However, the range of positions for representing texts or visual data tends to differ. Contemporary visual generation systems are typically trained with extremely long and descriptive captions [4, 66], despite the low latent resolution of visual data (e.g., a video of resolution $448\times 256\times 25$ becomes $56\times 32\times 7$ after a $8\times 8\times 4$ compression). To balance the two modalities, we propose scaling the 3D positions to ensure a balanced learning. Specifically, we empirically scale the latent 3D positions to the RGB space by multiplying the compression ratio, as shown in figure 5(a). This simple scaling operation, from another perspective, improves the RoPE resolution for visual tokens by slightly accelerating the rotation speed. Experiments in the experiment section demonstrate its efficacy, thereby showing the importance of balancing the two modalities from the RoPE perspective. We acknowledge, however, that given the autoregressive generation nature of videos, this scaling may not be the optimal solution. More advanced and sophisticated solutions are left for future work.
3.2 Autoregressive Discrete Diffusion Forcing
The most naive generation paradigm (i.e., next-token prediction) suffers from low generation efficiency, rendering it impractical for autoregressive visual generation. In this work, we resort to discrete diffusion [81, 63] to generate visual content, together with temporally causal dependency to enable a temporally autoregressive generation paradigm. However, due to the autoregressive nature of Lumos-1, naive random masking (i.e., a globally random mask) or temporally independent masking (i.e., diffusion forcing [9, 60]) both lead to significant loss imbalance, i.e., visual tokens in the later frames tend to have much lower loss. Since the task difficulty of predicting frames with ample history frame context is considerably easier than predicting the first image given a text prompt or predicting the second frame given the first frame, the model would lead towards optimizing simpler tasks, leading to a degradation of temporal learning.
Training scheme. To resolve this problem, we build upon the basic characteristics of videos–spatial information redundancy. The core problem of imbalanced loss during training stems from spatial information leakage. It is worth noting that the original video diffusion transformer [60] that adopts diffusion forcing does not encounter this issue due to its usage of bidirectional dependency. Facing this challenge, we introduce autoregressive discrete diffusion forcing (AR-DF), a technique that adopts temporal tube masking during the training of autoregressive video generation. Concretely, for every video, we randomly generate a mask pattern for the first frame and then apply this mask pattern repeatedly to later frames in this video. If we denote the multi-modal token sequence $\bm{X}$ composed of text tokens $\bm{X}_{p}$ and visual tokens $\bm{X}_{v}$ and sample a mask ratio $\rho$, the training masking strategy in AR-DF can be formulated as:
$\small\bm{M}_{i}\sim\mathrm{Bernoulli}(1-\rho)\quad\text{for }i=1,\dots,N_{f}$ (7)
$\widetilde{\bm{X}}_{v}^{(t)}\;=\;\bm{M}\odot\bm{X}_{v}^{(t)}\;+\;(1-\bm{M})\,\odot\,[\text{MASK}],\quad t=1,\dots,T$ (8)
$\small\widetilde{\bm{X}}\;=\;\bigl{\{}\bm{X}_{p},\;\widetilde{\bm{X}}_{v}^{(1)},\;\widetilde{\bm{X}}_{v}^{(2)},\;\dots,\;\widetilde{\bm{X}}_{v}^{(T)}\bigr{\}}$ (9)
[…]
Table 1: Performance comparison on GenEval. In the “#Params" column, we present the parameter counts of the visual generative model and the significant language encoders for a fair comparison with unified models following [95]. Tables below follow this paradigm.
[…]
Diffusion models
SD v1.5 [55] 0.9B + 0.1B 2B 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD v2.1 [55] 0.9B + 0.3B 2B 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL [48] 2.6B + 0.8B 0.55 0.98 0.74 0.39 0.85 0.15 0.23
SD 3 [15] 8.2B + 2.8B 0.68 0.98 0.84 0.66 0.74 0.40 0.43
DALL-E 2 [50] 4.2B + 1.0B 650M 0.52 0.94 0.66 0.49 0.77 0.10 0.19
FLUX [32] 12B + 2.5B 0.67 0.99 0.85 0.75 0.77 0.22 0.42
Autoregressive models
LlamaGen [62] 0.8B + 2.9B 60M 0.32 0.71 0.34 0.21 0.58 0.07 0.04
Show-o [81] 1.3B 2B 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Chameleon [65] 34B 1.4B 0.39
Transfusion [95] 7.3B 3.5B 0.63
EMU3 [95] 8B 0.66 0.99 0.81 0.42 0.80 0.49 0.45
LWM [39] 7B 1B 0.47 0.93 0.41 0.46 0.79 0.09 0.15
SEED-X [18] 17B 0.49 0.97 0.58 0.26 0.80 0.19 0.14
Lumos-1 ($352\times 352$) 1.5B 60M 0.601 0.959 0.732 0.375 0.774 0.365 0.400
Lumos-1 ($352\times 352$) 3.6B 60M 0.664 0.953 0.806 0.463 0.806 0.483 0.475
3.3 Implementation
Architecture. The architecture of Lumos-1 follows Llama [71, 20], meaning that it incorporates RMSNorm [92] and SwiGLU [57] by default. To stabilize training, we integrate query-key normalization (QK-Norm) following Chameleon [65]. We have models of three scales (Lumos-1 0.5B, 1B and 3B), whose architectural details are placed in the Appendix. It is worth noting that we use the 0.5B version for fast ablation studies.
Tokenizers. To unify visual and text token processing, we adopt the discrete version of Cosmos Tokenizer [1] that achieves spatiotemporal compression rates of $8\times 8\times 4$. For text tokens, we retain Chameleon’s text tokenizer [65]. Therefore, Lumos-1’s total codebook size is 129,536, partitioned into 65,536 text tokens and 64,000 visual tokens.
Sequence formatting. The visual tokens and text tokens are interleaved within a sequence [38], with the text tokens specifying metadata, including the text prompt, video resolution, video fps and the number of frames in this video. With this design, we train the model using images and videos of varying aspect ratios without resizing them. The details of the sequence formatting (e.g., organization of special tokens) are placed in the Appendix.
Table 2: Performance comparison on VBench-I2V benchmark. We list partial metrics due to space limits.
[…]
Diffusion models
VideoCrafter-I2V [11] 2.6B 20M 82.57 86.31 78.84 91.17 91.31 97.86 98.79 71.68
ConsistI2V [54] 1.3B + 0.3B 10M 84.07 91.91 76.22 95.82 95.95 95.27 98.28 66.92
SEINE [12] 0.9B + 0.1B 10M 84.88 92.39 77.37 96.57 96.80 94.2 97.26 70.97
I2VGen-XL [94] 1.4B + 1.0B 35M 85.28 92.11 78.44 96.48 96.83 94.18 97.09 69.14
CogVideoX [83] 5.6B + 4.8B 86.70 94.79 78.61 97.19 96.74 94.34 96.42 70.01
Autoregressive models
COSMOS-Video2World [62] 5B + 11B 100M 84.16 92.51 75.81 95.99 97.36 97.12 96.59 59.90
Lumos-1 ($672\times 384\times 25$) 1.5B 10M 84.16 91.80 76.53 96.06 96.58 95.90 96.25 67.06
Lumos-1 ($672\times 384\times 25$) 3.6B 10M 84.72 93.34 76.10 97.42 97.40 97.42 96.91 69.23
Table 3: Performance comparison on VBench-T2V benchmark. We list partial metrics due to space limits.
[…]
Diffusion models
ModelScopeT2V [73] 1.4B + 0.3B 10M 75.75 78.05 66.54 89.87 95.29 58.57 82.25 81.72 25.67
InstructVideo [91] 1.4B + 0.3B 10M 76.61 81.56 56.81 95.30 96.97 68.01 73.26 77.14 19.91
LaVie [78] 2.5B + 0.5B 25M 77.08 78.78 70.31 91.41 97.47 61.90 91.82 86.39 26.41
OpenSoraPlan V1.3 [35] 2.7B + 13B 70M 77.23 80.14 65.62 97.79 97.24 56.21 85.56 79.30 24.47
LTX-Video [22] 1.9B + 4.8B 80.00 82.30 70.79 96.56 97.20 60.28 83.45 81.45 25.19
CogVideoX [83] 5.6B + 4.8B 81.91 83.05 77.33 96.45 96.71 63.33 85.07 83.03 27.65
Autoregressive models
CogVideo [26] 9B 5.4M 67.01 72.06 46.83 92.19 96.20 41.03 73.40 79.57 7.70
EMU3 [77] 8B 80.96 84.09 68.43 95.32 97.69 86.17
Lumos-1 ($448\times 256\times 25$) 1.5B 10M 75.78 77.54 68.74 94.97 96.08 54.58 90.08 75.34 24.71
Lumos-1 ($448\times 256\times 25$) 3.6B 10M 78.32 79.52 73.51 95.51 96.50 58.04 90.05 82.00 25.29
[…]
Stage-wise training. Due to the autoregressive nature of Lumos-1, the training of video generation can be categorized into the training of two capabilities: 1) text-to-image and 2) image/images to video. Although the incorporation of AR-DF training substantially ameliorates the imbalanced learning issue, we still observe that the later task is relatively easier. Therefore, a stage-wise training scheme is mandatory to ensure successful video generation training. Concretely, the training commences with dedicated text-to-image training with 256p resolution, and then we perform image-video joint training to train the video generation capability on 256p resolution. Finally, we perform joint training on visual data of 384p resolution.
4 Experiments
4.1 Experimental Details
Datasets. To train Lumos-1, we curate a image dataset containing 60 million images and a video dataset containing 10 million videos. We preserve their original aspect ratios, with videos being clipped to 25 frames for training. To ensure fine-grained vision-language alignment [50, 66], the visual data is re-captioned using visual-language models [36] to obtain long and descriptive captions.
Training, inference and model evaluation. We train the model from scratch and the basic training hyper-parameters are placed in the Appendix. We adopt the Adam [41] optimizer to optimize the model, set the weight decay to 0.1 and set $\beta_{1}$ and $\beta_{2}$ to $0.9$ and $0.95$. During AR-DF training, the training mask ratio $\rho_{tra}$ is set to $0.7$ following [33]. During AR-DF inference, classier-free guidance (CFG) is utilized by default to enhance the generation quality. The inference mask ratio $\rho_{inf}$ is set to $0.7$ by default. The number of steps to generate one latent frame is set to $N_{steps}=50$ by default. When evaluating on GenEval [19] and VBench [28], the guidance scale is set to 16 if not otherwise specified. Since we train our model only on detailed and descriptive captions, it is mandatory for us to rewrite captions using Qwen 32B [82] when evaluating on GenEval [19]. When evaluating on VBench [28], we use its official long captions by default.
4.2 Comparison with Other Methods on Visual Generation
Text-to-image generation. We compare Lumos-1 with competitive image generation methods in Tab. 1. Compared with diffusion models, we observe that Lumos-1 outperforms models of similar sizes by a margin (e.g., SD-XL [48]) and is even on par with FLUX [32]. Compared with autoregressive models, we observe that Lumos-1 is on par with EMU3 [77], while being significantly more efficient due to the discrete diffusion inference. Lumos-1 achieves superior results in terms of position and attribute binding, demonstrating excellent language understanding and vision-language alignment even without textual pre-training.
Image-to-video generation. Thanks to the autoregressive nature of Lumos-1, we can perform image-to-video generation by specifying the first frame, although we did not specifically train on this task. Results are listed in Tab. 2. Lumos-1 outperforms the popular VideoCrafter-I2V model and is on par with the leading COSMOS-Video2World model, which uses substantially more data (100M$>$10M) and training resources (10000 H100s$>$48 H20s), demonstrating the promising performance of Lumos-1.
Text-to-video generation. We list the comparison in Tab. 3. Although we utilize discrete tokenizers, Lumos-1 can still be on par with diffusion models like OpenSoraPlan even if we do not rely on a heavy pretrained text encoder. Due to the autoregressive generation nature, the video quality can be ensured by first frame quality, enabling Lumos-1 to excel in object-centric metrics (object class and color).
4.3 Analysis and Ablation Studies
Figure 6: Visual comparison of Lumos-1 with other methods on text/image-to-video tasks.
Qualitative visual comparison. We compare Lumos-1 with popular video generation methods in figure 6. For T2V, the visual quality of our 384p videos does not trail 512p videos from LTX-Video. In the provided case, Lumos-1 generates better natural motion (water waves) and aligns with prompts better (skier in red and waves). For I2V, Lumos-1 handles multiple-object (multiple floating hot air balloons in example 1) and fine-grained motion (subtle ripples around the shoreline in example 3) substantially better than Stable Video Diffusion [5] (SVD), which generates global camera movement only. In example 2, SVD produces significant blurs, whereas Lumos-1 animates subjects smoothly. More visualizations are placed in the Appendix.
Effectiveness of temporal-tube masks during AR-DF training. In figure 7(a), we compare the frame-wise validation loss (frame 0, 3, 6) when using global random masks and temporal tube masks. The training mask ratios $\rho_{tra}$ are both set to $0.7$ following MAR [33]. For random masks, the loss of frame 6 decreases immediately and becomes lower than that of earlier frames. This rapid fall indicates pronounced information leakage: the model can reconstruct a masked token by attending to unmasked tokens in neighbouring frames instead of modelling genuine temporal dynamics, rendering the task overly easy. For temporal tube masks, frame 6 is the hardest because all pixels in the same spatial locations are masked across the temporal axis, eliminating the shortcut available with random masks. As iterations proceed, the gap between frames narrows and eventually levels out, demonstrating that the model is learning to propagate information through time rather than copying it.
Figure 7: (a) Effectiveness of temporal-tube masks during AR-DF training (on videos) using the 0.5B model; (b) Effectiveness of MM-RoPE; (c) Sensitivity analysis of CFG scale on GenEval using 1B model.
Figure 8: (a) Selected VBench metrics with varying inference-time mask ratio $\rho_{inf}$; (b) Text-to-video visualization comparing the effect of inference-time masks. (w/o: $\rho_{inf}=0.0$, w/: $\rho_{inf}=0.7$)
Effect of AR-DF inference masks and sensitivity to $\rho_{inf}$. AR-DF requires the same partial-context masking at inference as during training; omitting these masks severely harms quality. In figure 8(b), we observe that the “w/o inference mask" setting produces visible artifacts and flickering, whereas using masks preserves coherence. In figure 8(a), we select two metrics (imaging quality and dynamic degree) from VBench to quantitatively assess the impact of $\rho_{inf}$. A broad plateau between $0.3$ and $0.7$ yields smooth, visually pleasing videos. When $\rho_{inf}$ is below $0.3$, insufficient context forces the model to degrade both motion and per-frame quality, thus inflating dynamic degree values. When $\rho_{inf}$ is above $0.7$, overly aggressive masking disrupts temporal continuity, driving up dynamic degree values. We empirically set $\rho_{inf}$ to $0.7$ to ensure obvious motion.
Table 4: Inference time with different RoPEs using 1B / 3B model. Results are reported on one H20 with batch size set to 1, $N_{steps}=50$ and CFG used by default.
1D RoPE M-RoPE MM-RoPE
Image ($448\times 256$) 7.4s / 16.3s 7.7s / 16.9s 7.7s / 16.9s
Video ($448\times 256\times 25$) 75.1s / 173.6s 77.8s / 178.5s 77.8s / 178.5s
Effectiveness of MM-RoPE. figure 7(b) plots the validation loss of the 0.5B model under four RoPE settings. Note that M-RoPE means that both designs are removed. We can observe that MM-RoPE consistently converges faster and settles at the lowest loss, confirming the benefit of modelling fine-grained spatiotemporal information. Although ablating either component raises the loss, dropping the distributed design hurts more than dropping the design of scaled position, indicating that comprehensive frequency allocation is the dominant factor. Removing both enhancements gives the slowest convergence and the highest plateau, showing that the two mechanisms are complementary for effective video generation.
Figure 9: (a) Effect of the number of meta MM-RoPEs in MM-RoPE; (b) Effect of the scaling factors in MM-RoPE. Both curves are plotted using the 0.5B model trained on videos.
Effect of the number of meta MM-RoPEs in MM-RoPE. MM-RoPE slices the embedding channels into several meta groups. More groups mean that one kind of information (temporal, height or width) receives a wider and more comprehensive frequency spectrum for modeling. figure 9(a) plots the validation loss for a 0.5B model under four settings: 1) w/o distributed design: This setting uses the previous design, which allocates the first $\frac{2}{8}$ channels for temporal modeling and $\frac{3}{8}$ channels for height and width modeling, respectively. 2) #meta MM-RoPE=1: This setting equips a meta MM-RoPE of 64 channels, while maintaining the ratios ($2:3:3$) for modeling temporal, height and width information. This variant improves upon the previous one by enabling interleaved height and width channels, which increases the frequency spectrum range for the two spatial dimensions. 3) #meta MM-RoPE=2: This setting equips a meta MM-RoPE of 32 channels. This variant improves the frequency spectrum range for temporal, height and width information compared with the previous one. 4) #meta MM-RoPE=4: This setting is the default design that keeps the number of channels for every meta MM-RoPE minimal (16 channels). This means that the frequency spectrum range is most comprehensive for the temporal, height, or width dimension. These results confirm that widening the frequency spectrum for every dimension through increasing the number of meta MM-RoPEs substantially improves spatiotemporal modeling and overall training effectiveness.
Effect of the scaling factors in MM-RoPE. figure 9(b) presents the validation loss curves of varying the scaling factors for modeling temporal, height and width positional information in MM-RoPE. Two clear trends emerge: Moving from $(1,1,1)$$(2,4,4)$$(4,8,8)$ steadily lowers the curve, but further enlarging to $(8,16,16)$ or $(16,32,32)$ yields no additional gain since the three curves almost overlap throughout training. Therefore, a moderate scale of $(4,8,8)$ is sufficient to balance vision-language ranges and fully realize the benefit of higher-resolution RoPE, while avoiding unnecessary frequency inflation. We therefore adopt $(4,8,8)$ as the default scaling for MM-RoPE.
Table 5: GenEval score across different aspect ratios.
$\bm{448\times 256}$ ($7:4$) $\bm{352\times 352}$ ($1:1$) $\bm{256\times 448}$ ($4:7$)
GenEval 0.605 0.601 0.569
Inference overhead analysis for MM-RoPE. Similar to M-RoPE, MM-RoPE needs to locate the start of visual tokens and then apply the RoPE mechanism, which requires a small amount of computation. In Tab. 4, we compare the inference speed of generating images and videos using the vanilla 1D RoPE, M-RoPE and MM-RoPE. We can observe that: 1) Compared with 1D RoPE, the incorporation of 3D priors introduces only 3.5%-4.1% inference latency. 2) Compared to M-RoPE, MM-RoPE introduces no additional latency.
Sensitivity analysis of CFG scale. We study the impact of the guidance scale on GenEval using 1B model in figure 7(c). We find that scale values from 13 to 16 (default value) lead to decent results.
Robustness to aspect ratios. Although the aspect ratio of training data is mostly $7:4$, Tab. 5 indicates that Lumos-1 1B adapts well to visual generation with different aspect ratios due to the unified codebook design.
5 Conclusion
In this paper, we introduce Lumos-1, a model that utilizes the LLM architecture for autoregressive video generation. We propose MM-RoPE for better spatiotemporal dynamics modeling and propose AR-DF for effective training and inference considering intra-frame bidirectionality and inter-frame temporal causality. We anticipate that Lumos-1 represents a significant step toward building a foundational unified model.
Acknowledgement
The LaTeX template is built upon Meta’s original template.
Appendix
In this Appendix, we provide additional content organized as follows:
  • appendix A presents a detailed version of the AR-DF inference algorithm.
  • appendix B presents more details concerning architectures and implementation.
  • appendix C provides more visualizations to demonstrate image and video generation capabilities.
  • appendix D presents discussions covering 1) limitations and future work (Sec. D.1) and 2) potential societal impact and safeguards (Sec. D.2).
Appendix A Detailed AR-DF Inference Algorithm
In this subsection, we present the detailed version of the AR-DF inference algorithm. After the text prompt $\bm{X}_{p}$ is encoded once and its KV pairs are cached (Lines 1–3), we iterate over the $T$ latent frames (Line 6). Each frame is generated by running $N_{steps}$ iterations (Lines 9–22). During the frame generation process, the model predicts a token distribution, samples tokens only at positions that are still masked, gathers their confidences, and sets the scores of already-revealed tokens to +$\infty$+ ∞. After adding Gumbel noise, we re-mask the lowest-confidence tokens. As the generation proceeds, fewer tokens are re-masked as the running steps grow and the frame becomes sharper and clearer. The partially revealed frame is then added to the cache before the next frame is generated. Once all frames are finished, they are decoded to RGB space to form the final video.
Algorithm 3 AR-DF Inference Procedure
Require: text prompt $\bm{X}_{p}$, trained model $G_{\phi}$, inference mask ratio $\rho_{inf}$, number of latent frames $T$, number of generation steps $N_{steps}$, number of tokens in a latent frame $N_{f}$, KV cache $\mathcal{C}=\varnothing$, generated latent list $\mathcal{V}=\varnothing$
1:Initialize text cache:
2: Generate text causal mask $AttnMask^{(p)}$ for $\bm{X}_{p}$
3:$\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$triangleright Store cache for the prompt
4:Sample cache mask:
5:$\bm{M}_{inf}\sim\mathrm{Bernoulli}(1-\rho_{inf})$
6:for $t=1\to T$ do $\triangleright$ Reused for all frames
[…]
8:    Generate temporal causal mask $AttnMask^{(t)}$ for $\{\bm{X}_{p},\bm{X}_{v}^{(t)}\}$
9:    for $n=1\to N_{steps}$ do
10:         $\bm{P}\leftarrow G_{\phi}\bigl{(}\bm{X}_{v}^{(t)},AttnMask^{(t)},\mathcal{C}\bigr{)}$triangleright Predicted token distribution for all tokens
[…]
15:         Mask the lowest confident ($\bm{Q}$) tokens in sampled_ids with randomness:
16:          Perturb $\bm{Q}$ with gumbel noise
17:          $\alpha\leftarrow\cos(\frac{\pi}{2}\frac{n}{N_{steps}})$triangleright Cosine schedule factor in [0,1]
18:          $U\leftarrow\text{floor}(\alpha\times N_{f})$triangleright Number of tokens to mask in this iteration
[…]
22:    end for
23:    Append $\bm{X}_{v}^{(t)}$ to $\mathcal{V}$
24:    Cache partial observation of the generated frame:
25:     $\widetilde{\bm{X}}_{v}^{(t)}=\bm{M}_{inf}\odot\bm{X}_{v}^{(t)}+(1-\bm{M}_{inf})\odot[\text{MASK}]$
26:     $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$triangleright Store cache for the $t$-th frame
27:end for
28:Decode $T$ latents $\mathcal{V}$ to RGB space
29:return the generated video frames
Figure A1: Detailed inference algorithm with AR-DF.
Appendix B More Architectural and Implementation Details
Model architectures and more training details. Tab. B1 presents the architectural details of three versions of Lumos-1 and their corresponding training details, including their training batch sizes and their learning rates. For 256p joint training, we use all 60M images and 10M videos. For 384p joint training, we curate a higher-quality subset from 256p training data, consisting of 8M images and 0.6M videos. Following the alternating strategy of MovieGen [49], we interleave image and video batches during joint training.
[…]
Table B1: Model architectural and training details. Batch sizes (*/*) are for images and videos during joint training. Learning rates (*/*) are for 256p and 384p. The 0.5B version is only for fast ablation studies.
Model #Params #Layers Hidden size #Heads Head dim Batch (256p) Batch (384p) Learning rate
Lumos-1 0.5B 0.5B 16 1024 16 64 896 / 128 192 / 32 5e-5 / 2.5e-5
Lumos-1 1B 1.5B 16 2048 32 64 640 / 96 192 / 32 5e-5 / 2.5e-5
Lumos-1 3B 3.6B 28 3072 24 128 768 / 96 288 / 48 1e-4 / 2.5e-5
Appendix C More Visualizations
Qualitative visual comparison on text-to-image generation. We compare Lumos-1 with popular text-to-image generation methods in figure C1. All models are configured to their default inference settings and use the same version of detailed prompts to ensure quality generation. We can observe that: 1) Although EMU3 is significantly larger than Lumos-1, Lumos-1 generates images that reproduces fine prompt details more faithfully, such as the dappled shadows on the bench (example 1) and the correct “STOP” text and the yellow color on the road sign (example 3). EMU-3 sometimes drops or warps these details and exhibits occasional anatomical distortions, such as the distorted zebra (example 5) and malformed fingers on the woman holding an apple (example 7). 2) Stable Diffusion 3 offers slightly crisper textures, which is expected from its continuous tokenizer, but it frequently ignores descriptive constraints, such as the dappled shadows (example 1) and gentle shadows (example 4), and the incorrect dress color (example 7). By contrast, Lumos-1 meets these constraints while maintaining a convincing overall composition. These examples illustrate that Lumos-1 delivers strong vision–language alignment compared with both diffusion and autoregressive baselines, while remaining competitive in visual quality.
Figure C1: Visual comparison of Lumos-1 with EMU3 and Stabel Diffusion 3 (Medium) on the text-to-image task. We place the simple version of the prompts above the images for reference. Fragments of the detailed prompt are placed in parentheses, with the critical attributes highlighted in red. All images are generated using the detailed prompts for all models.
More text-to-image visualizations. figure C2 presents more results generated by Lumos-1 using descriptive captions. Due to the unified codebook, we can train Lumos-1 using the original resolution of the visual data, enabling generation with different aspect ratios such as $4:7$, $7:4$, and $1:1$, although we do not have a significant amount of data with aspect ratio $4:7$. Even though Lumos-1 is trained from scratch, it transforms lengthy descriptions into complex scenes that are aligned faithfully with the prompts. This can be exemplified by the second row of images, where these images clearly reflect the prompts:
  • a photo of three handbags” contains the description: To the left, a classic black leather handbag with a structured design and gold-tone hardware catches the eye. In the center, a vibrant red quilted handbag with a chain strap adds a pop of color and a touch of luxury. On the right, a casual beige canvas tote bag with simple, clean lines offers a more relaxed aesthetic.
  • a photo of four vases” contains the description: The first vase on the left is a tall, cylindrical glass vase with a slight teal tint, showcasing clear transparency and a modern design. Next to it is a vintage ceramic vase with an off-white base adorned with hand-painted blue floral patterns, adding a touch of traditional elegance. The third vase is a short, squat terracotta pot with a rustic, earthy texture and visible firing marks, evoking a natural, bohemian feel. Finally, a sleek, metallic silver vase with a geometric shape completes the set, providing a contemporary contrast.
  • a photo of four dogs” contains the description: To the left is a golden retriever, its fur glowing warmly in the sunlight, with a gentle, friendly gaze. Next to it is a sleek black Labrador, its coat shining with health, head tilted slightly as if listening intently. In the center, a small, energetic Jack Russell terrier bounces with excitement, its white and brown fur clearly defined. On the right, a large German Shepherd sits calmly, its regal posture and dark, expressive eyes adding a dignified presence to the group.
These qualitative results underscore the strong vision–language alignment achieved by Lumos-1.
Figure C2: More text-to-image results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolutions for three rows of images are $384\times 672$, $672\times 384$ and $512\times 512$.
More image-to-video and text-to-video visualizations. figure C3 presents more image-to-video results and figure C4 presents more text-to-video results generated by Lumos-1. The prompts span a wide semantic and dynamical range, containing 1) close-up animal/human scenes (example 4, 5 in figure C4), 2) fast human and vehicle motion (example 1, 4 in figure C3 and example 1, 2 in figure C4), 3) fluid dynamics (example 1, 5 in figure C3 and example 2 in figure C4), and 4) complex multi-object collective motion (example 3 in figure C3 and example 3 in figure C4). The results demonstrate that Lumos-1 can 1) handle videos of different aspect ratios and 2) generate single-subject and multi-subject motion that is temporally coherent and physically plausible.
Figure C3: More image-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The results encompass videos of resolution $672\times 384$ and $768\times 320$.
Figure C4: More text-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolution is $672\times 384$.
Appendix D Discussions
D.1 Limitations and Future Work Discussions
We recognize that Lumos-1, as an initial endeavor in this area, comes with its limitations. Most prominently, its training corpus (60 million images and 10 million videos) is modest compared with datasets used by recent foundation models [66, 32], which usually contain billions of samples. Consequently, Lumos-1 can under-generalize in scenarios that require fine-grained human actions or highly intricate scene dynamics. Considering this, our immediate research plan therefore includes scaling along three axes: 1) Data volume and diversity – expanding both image and video coverage to narrow the generalisation gap. 2) Model capacity – training larger backbones while retaining the MM-RoPE and AR-DF designs. 3) Multimodal knowledge infusion – initializing with strong vision–language models or co-training with visual understanding tasks so the generator can better ground its outputs in world knowledge.
D.2 Potential Societal Impacts and Safeguards
As the pioneering efforts in autoregressive visual generation, Lumos-1 can be a significant step towards a large-scale unified model for general visual understanding and generation. Through substantially large-scale training, the obtained model could stimulate intelligence that exceeds existing models trained on unimodal data. The obtained model could serve as a foundation for various downstream applications, thereby reducing the carbon and economic cost that society spends on training various specialized models.
At the same time, unrestricted deployment carries non-trivial risks: the model might be used to fabricate deceptive media, produce disallowed or disturbing visuals, or amplify harmful biases present in the training data. To mitigate these concerns, we recommend the following safeguards before any real-world release: 1) Alignment tuning - apply preference- or instruction-tuning so that outputs respect human aesthetic and moral preferences [91]. 2) Dual-stage filtering – i) prompt filters that refuse generation requests involving illicit content, and ii) post-generation classifiers that block or watermark unsafe outputs. 3) Red-team evaluation – continuous adversarial testing to uncover failure modes in new domains or under distribution shift. Lumos-1 is a research-oriented work, and any broader deployment should proceed only with thorough oversight and evaluation to ensure responsible and ethical use.